OpenStudy (briensmarandache):

use addition formula to simplify the equation. then find all real solutions: sin(2x)cos(x)+cos(2x)sin(x)=1/2(sqrt(3))....... im not seeing how to apply the addition formula

OpenStudy (campbell_st):

the left hand side is in the form sin(A + B) = sin(A)cos(B) + cos(A)sin(B) hope that is the start you need...

OpenStudy (briensmarandache):

could you expand on the rest of the problem. not seeing how im supposed to find solutions @SolomonZelman i got the top part

OpenStudy (solomonzelman):

re-write the left side of the problem, using the rule that campbell posted.

OpenStudy (briensmarandache):

sin(2x+x)=.5(sqrt(3))

OpenStudy (briensmarandache):

would i just take the inverse of sine and divide by 3?

OpenStudy (briensmarandache):

i dont know if that is algebraically appropriate

OpenStudy (briensmarandache):

@campbell_st so is that right?

OpenStudy (briensmarandache):

@SolomonZelman